8. An interior decorator is buying wallpaper for two square walls. One wall
is 3 feet wider than the other. If he needs 89 square feet of wallpaper
, what are the dimensions of the wall?

Answer by ikleyn(52786)

.
An interior decorator is buying wallpaper for two square walls. One wall
is 3 feet wider than the other. If he needs 89 square feet of wallpaper,
what are the dimensions of the wall?
~~~~~~~~~~~~~~~~~

x^2 + (x+3)^2 = 89


x^2 + (x^2 + 6x + 9) = 89


2x^2 + 6x - 80 = 0


x^2 + 3x - 40 = 0


(x+8)*(x-5) = 0


Of the two roots, -8 and 5, only positive root  x= 5 is the solution.


Answer.  One wall is 5 ft square; the other wall is 8 ft square.

Solved.

-------------

The question in the post is posed out of the English standard rules and out the human logic - which just becomes a TRADITION at this forum.